# A Closer Look at Knowledge Distillation in Spiking Neural Network Training

### v1

@@ Introduction
Spiking Neural Networks (SNNs), inspired by the spiking mechanism of biological neurons, utilize event-driven binary spikes to transmit information, allowing multiplications between activations and weights to be replaced by additions or remain silent, thereby significantly improving energy efficiency (eshraghian2021training; lecun2015deep; davies2018loihi). Taking advantage of this computational paradigm, SNNs can operate efficiently on neuromorphic hardware and demonstrate autonomous learning capabilities and ultralow power consumption (roy2019towards; mehonic2022brain; schuman2022opportunities), making them highly promising for intelligent computing tasks (fang2023spikingjelly; zhang2020system). However, training SNNs presents significant challenges due to the inherently discrete and sparse nature of spike-based features, which complicates their optimization process and results in performance and application limitations compared to traditional artificial neural networks (ANNs) (zhou2024towards; zhou2024label; zhou2022audio; zhou2024avss; zhou2024dense; zhou2021positive; zhou2024vaplan; zhou2023contrastive; zhang2025temporal; li2025patch; zhao2025multimodal).
@@
Converting a pre-trained ANNs to SNNs (ANN-to-SNN) and directly training are two common approaches for obtaining SNNs. Specifically, conversion-based methods (bu2022optimal) transfer pre-trained ANN parameters to corresponding SNNs but replace ReLU activation with spiking neurons (huang2024clif). This strategy has been experimentally found to require a large number of time steps to achieve satisfactory performance (bu2022optimal; han2020rmpsnn; Rueckauer2017conversion). Direct training methods (wu2019direct; fang2021sew; deng2022tet), on the other hand, optimize SNNs using direct backpropagation through the surrogate gradient estimation technique (zenke2021remarkable), leading to significant progress in pattern recognition (deng2022tet; zhou2023spikformer; deng2024spikingtoken), natural language processing  (zhang2024spikingminilm; xiao2022towards), and multimodal tasks (LiuMISNet2025). While this strategy reduces training time steps, the SNN’s performance still lags behind that of ANNs.
@@
Recent works (xu2025bkdsnn; Xu2023kdsnn) improve SNN training using knowledge distillation (KD) (hinton2014distilling) techniques, with pretrained ANNs as the teacher model and SNNs as the student model, showing promising results on multiple datasets. These methods perform KD by element-wise aligning the output features (Xu2023kdsnn) or classification logits (Xu2023kdsnn) between ANNs and SNNs. Such a KD paradigm distills knowledge from teacher ANNs to student SNNs, however, prior works ignore two critical issues: (i) Discrepancy between raw features. As illustrated in Fig. 1(a), features extracted from ANNs via a single forward pass are represented as continuous floating-point values. In contrast, features in SNNs, obtained by forward propagation over multiple time steps, are expressed as discrete binary spikes. Moreover, whereas ANN features encapsulate patterns spanning the entire image, SNN spikes primarily highlight salient regions. (ii) Discrepancy between raw logits. The logits (i.e., raw classification scores) are derived from the hidden features. Consequently, a notable disparity emerges between the raw logits of the two models. Specifically, SNN logits display greater sparsity and a more peaked distribution relative to those of ANNs.
@@
In this paper, we propose novel knowledge distillation strategies, to address the above challenges. Specifically, to address the first challenge, we propose the Saliency-scaled Activation Map Distillation (SAMD). As illustrated in Fig. 1, unlike previous methods that directly perform knowledge distillation using raw features, our SAMD leverages the Class Activation Map (CAM) (zhou2016learning) of the teacher ANN, which provides more precise and focused knowledge, clearly describing the salient image regions related to the target class. Notably, unlike traditional activation map-based distillation methods from ANNs (e.g., e2KD (parchami2024good) and CATKD (guo2023class)), we discover that the surrogate gradient estimation in SNNs prevents the use of precise gradient estimation methods like Grad-CAM (selvaraju2017grad) to generate saliency activation maps. Instead, we redesign the activation map distillation by aligning the Spiking Activation Map (SAM) of the student SNN with the CAM of the teacher ANN. Although both SAM and CAM originate from features, they are more consistent than raw features due to the use of saliency maps. Furthermore, we consider the numerical magnitude differences between the activation maps generated from SNN features (i.e., SAM) and ANN features (i.e., CAM). To more accurately assess the contribution of each pixel in the saliency maps, we apply the softmax function to convert both CAM and SAM into probability distributions. In this way, the scaled CAM and SAM remain consistent in both semantics and numerical magnitude, facilitating their alignment. To address the second challenge, we propose the Noise-smoothed Logits Distillation (NLD). As demonstrated in Fig. 1(b), NLD employs Gaussian noise to moderate the prediction logits of student SNNs. Specifically, we sample Gaussian noise with mean and variance parameters derived from the SNN logits, ensuring that the original distribution of these logits remains largely preserved. After the addition of noise, the logits of SNNs transit from a sparse and sharply peaked distribution to one that is denser and broader, resembling the distribution of ANN logits, thereby facilitating knowledge transmission between teacher and student. We evaluate the effectiveness and superiority of the proposed two KD strategies on CIFAR-10, CIFAR-100, and ImageNet-1K da of SNNs.
@@
Spiking Neural Networks (SNNs) are brain-inspired models that mimic biological neural systems by transmitting information through discrete spikes, achieving lower energy consumption. SNNs are typically trained using two main approaches: ANN-SNN conversion (meng2022training; bu2022optimal; deng2021optimal; hu2023spiking) and direct training  (fang2021sew; guo2023rmp; guo2023mbpn; deng2022tet; meng2023towards). The conversion methods directly transform pre-trained ANN into SNN by replacing its ReLU activation functions with integrate-and-fire (IF) (bu2022optimal) neurons. However, the converted SNNs often require prolonged time steps to collect sufficient spike signals to ensure accuracy. The direct training methods directly train the SNN by backpropagating surrogate gradients (fang2021sew) through multiple time steps, which can alleviate excessive time steps, enabling efficient training and inference. However, a significant accuracy gap persists between the trained SNNs and ANNs due to the approximation errors (Xu2023kdsnn) inherent in surrogate gradients. Unlike them, our method utilizes the informative knowledge in pretrained ANNs to better supervise SNN training, achieving balance between accuracy and efficiency.
@@
Knowledge Distillation (KD) (hinton2014distilling) was initially proposed for compressing artificial neural networks (ANNs) by transferring knowledge from complex teacher models to lightweight student models. Traditional logits-based KD methods typically minimize the difference in classification probabilities (i.e., logits) between the student and teacher models (sun2024logit; jin2023multi). In addition, feature-based KD methods enhance the performance of the student model by learning the feature representations of the teacher model (wang2025improving; guo2023class). Notably, a special class of feature-based KD methods (guo2023class; parchami2024good; zagoruyko2017paying) enhances the performance of the student by minimizing the semantic saliency activation maps of both the teacher and student models (e.g., Grad-CAM (selvaraju2017grad)). Recently, a few works start to apply KD technique to facilitate SNN training (xu2025bkdsnn; Xu2023kdsnn), significantly improving the model’s performance. i.e., KDSNN (Xu2023kdsnn) regularizes consistency of the output features and logits between ANNs and SNNs, while BKDSNN (xu2025bkdsnn) enhances feature-level matching by further processing the spike features of SNNs with a blurring matrix. However, these methods largely overlook the discrepancies of the raw features and logits from teacher ANNs and student SNNs. Specifically, ANNs generate continuous floating-point features, while SNNs generate discrete spike features. The logits of SNNs are also more sparse than that of ANNs. Thus, we perform KD from the perspective of semantic saliency activation maps, which are more semantically aligned. But, prior activation map-based KD methods (parchami2024good; zagoruyko2017paying) from ANNs cannot be directly applied to SNNs, due to the surrogate gradient estimation errors in SNNs, which affect the generation of semantic saliency activation maps. Besides, CATKD (guo2023class) is only applicable to CNN-based architectures, which limits the application of semantic saliency activation map-based KD in SNNs. In contrast, we design a semantic saliency activation map-based KD method for the characteristics of SNNs, which can be applied to various architectures. Moreover, we also design a noise-smoothing distillation method at the logits-level to further enhance the performance of SNNs.
@@
The overall pipeline of our CKDSNN framework is illustrated in Fig. 2. Given a pretrained ANN teacher model and a learnable SNN student model, the proposed CKDSNN aims to train the student model by effectively distilling knowledge from the teacher model from two aspects: 1) Saliency-scaled Activation Map Distillation (SAMD). The class activation map (selvaraju2017grad) obtained from the teacher ANN is used to guide the student SNN to fire spikes in salient image regions. During the distillation process, we address the magnitude discrepancy between the two types of activation maps by scaling them into the same range. 2) Noise-smoothed Logits Distillation (NLD). The classification output logits of the student model is smoothed using additional Gaussian noise, making the vanilla sparse logits distribution of student SNN close to the continuous logits distribution of teacher ANN. This facilitates more precise logit-level knowledge distillation. We first have a brief introduction to the spiking neuron used in SNN, which leads to the discrepancy on the features and logits of ANN and SNN. Then, we elaborate on the proposed SAMD and NLD strategies, respectively.
@@
Typical SNNs (fang2021sew; deng2022tet; meng2023towards; jiang2024tab; deng2024spikingtoken) adopt the integrate-and-fire (IF) neuron as the fundamental unit. Specifically, the IF neuron first integrates input currents by updating its membrane potential, and then compares it with a pre-set threshold to generate a spike signal, followed by a reset mechanism of the membrane potential. This process can be formulated as:
@@
where $I[t]$ is the input current at time step $t$ and V[t1]V[t{-}1] is the membrane potential at previous t1t{-}1 time step. $\gamma$ denotes the membrane time constant and $\Theta(\cdot)$ represents the Heaviside function (huang2024clif). The IF accumulates input currents to update the membrane potential $H[t]$. When $H[t]$ exceeds the threshold $V_{\text{th}}$, a spike $S[t]$ is generated, and the membrane potential is reset to $V_{\text{reset}}$ (meng2023towards).
In image classification, the above is attached to each feature encoding block of SNN. Specifically, each encoding block contains multiple linear transformation layers, batch normalization layers, and IF spiking neurons, as shown in Fig. 2(a). This generates discrete spike features $\boldsymbol{F}^{st}{\in}\mathbb{R}^{T\times C\times H\times W}$ over $T$ time steps, where $C$, $H$, and $W$ denote the number of channels, height, and width of the feature, respectively. The discrete spike features $\boldsymbol{F}^{st}$ generated by the IF activation mechanism differ significantly from the continuous features $\boldsymbol{F}^{te}{\in}\mathbb{R}^{C\times H\times W}$ produced by ANN. This leads to a significant discrepancy in the feature distribution between $\boldsymbol{F}^{st}$ and $\boldsymbol{F}^{te}$, resulting in different prediction logits. Prior KD methods (Xu2023kdsnn) directly match these two types of features or logits element-wise, ignoring the essential differences in feature representation and distribution, leading to suboptimal distillation results. We address these issues through carefully designed KD strategies.
@@
We propose the Saliency-scaled Activation Map Distillation (SAMD) to train SNN rather than distilling the element-wise discrepant features used in existing works (Xu2023kdsnn; xu2025bkdsnn). Specifically, SAMD consists of three steps:
1) Class Activation Map (CAM) Generation. The class activation map is initially used as a visualization tool to enhance the explainability of convolutional neural network (zhou2016learning), which can highlight the related image region of a specific class. In our work, we use it as the teacher’s knowledge for distillation. We follow the typical method of Grad-CAM (selvaraju2017grad) to generate the class activation map. Specifically, given an input image $x$, we first extract the intermediate features using a pretrained teacher ANN model, denoted as $\boldsymbol{F}^{te}{\in}\mathbb{R}^{C\times H\times W}$. Then, the gradients of $\boldsymbol{F}^{te}$ are calculated according to the forward loss between prediction and the ground truth label $y{\in}\mathbb{R}^{K}$ ($K$ is the total number of classes). The class activation map $\boldsymbol{M}^{te}{\in}\mathbb{R}^{H\times W}$ can be obtained by associating $\boldsymbol{F}^{te}$ with the class label $y$. This process can be formulated as follows:
@@
In addition to the feature-driven activation map distillation, we consider the target-driven logits distillation, which aims to align the output probability logits of teacher ANN and student SNN. Although prior work (Xu2023kdsnn) considered this, they largely overlook the discrepancy between the logits. This issue still originates from the utilization of different features in logits production where SNNs employ binary spike features that are either 0 or 1, whereas ANNs utilize floating-point features. Consequently, the logits of SNN are very sparse, while those of ANN are more dense and continuous. We propose the Noise-Smoothed Logits Distillation (NLD) to address this problem. Thus, we try to soften the logits of student SNN, $z^{st}{\in}\mathbb{R}^{1\times K}$, reducing the discrepancy from $z^{te}{\in}\mathbb{R}^{1\times K}$. Specifically, we add some continuous noise $\epsilon{\in}\mathbb{R}^{1\times K}$ onto $z^{st}$. To avoid destroying the original distribution of $z^{st}$ while maintaining the characteristic of classification logits, we opt to Gaussian noise with the mean and standard deviation of $z^{st}$:
@@
Datasets. We evaluate our method on three widely used image classification datasets and a neuromorphic dataset in this research area: CIFAR-10/100 (krizhevsky2009learning), ImageNet-1K (deng2009imagenet) and CIFAR10-DVS (li2017cifar10). The details of these datasets are provided in the supplementary materials.
Model Configuration. To ensure fair comparison, we determine the configurations of teacher ANN and student SNN models based on prior studies (guo2024enof; xu2025bkdsnn). Specifically, for the CIFAR-10/100 datasets, we use the ResNet-19/20 versions provided by TET (deng2022tet) and QCFS (li2023seenn), with their corresponding ANN versions as the teacher models; and we also test a ViT-based architecture, using ViT-S as the teacher ANN and Spikformer-4-384 (zhou2023spikformer) as the student model. For the ImageNet-1K dataset, we use the ResNet-18 and ResNet-34 pre-trained on ImageNet-1K as teacher models, while SEW-ResNet (fang2021sew) serves as the student model. For the CIFAR10-DVS dataset, we use ResNet-19 as the student, with its ANN architecture serving as the teacher, trained in the same way as EnOF (guo2024enof).
Implementation Details. To ensure consistency with prior studies, our implementation on CIFAR-10/100 and ImageNet strictly follows the established distillation architecture (guo2024enof; xu2025bkdsnn). Specifically, we set the hyperparameters as follows: $\mathcal{T}$ in Eq. Saliency-scaled Activation Map Distillation and $\tau$ in Eq. 7 to $2.0$. $\lambda$ in Eq. 6 to $0.1$, $\beta$ and $\gamma$ in Eq. 8 to $1.0$. The Integrate-and-Fire (IF) neuron settings align with prior works (fang2021sew), and all other training configurations, including batch size, learning rate, and optimizer, remain consistent with those in (xu2025bkdsnn). In addition, we conduct sensitivity analysis of all hyper-parameters in the supplementary materials.
Platform. All experiments are conducted on a server platform equipped with 32 cores Intel Xeon Platinum 8352V CPU with 2.10GHz and 8-way NVIDIA GPUs. We use SpikingJelly (fang2023spikingjelly) to simulate the IF (tal1997computing) spiking neurons.
@@
SEW-R (fang2021sew) NIPS’21 4 89.07 60.16 93.24 70.84 - -
STBP (wu2019direct) AAAI’21 4 - - 92.92 - - -
TET (deng2022tet) ICLR’22 4 - - 94.44 74.47 - -
SLTT (meng2023towards) ICCV’23 4 - - 94.56 74.67 - -
Spikformer (zhou2023spikformer) ICLR’22 4 - - - - 95.93 79.65
@@
KDSNN (Xu2023kdsnn) CVPR’23 4 89.03 60.18 94.36 74.08 95.88 80.33
BKDSNN (xu2025bkdsnn) ECCV’24 4 89.29 60.92 94.64 74.95 96.06 81.26
EnOFSNN (guo2024enof) NIPS’24 1 92.66 70.38 95.37 77.08 - -
@@
Results on CIFAR-10/100. The comparison results between our CKDSNN and previous methods are shown in Tab. 1. We significantly outperform existing methods across different architectures. For example, in the ResNet-19, when the time step is set to 1, CKDSNN achieves an accuracy improvement of 0.74% on the CIFAR-10 dataset and 1.03% on the CIFAR-100 dataset compared to the current most competitive knowledge distillation method, EnOF (guo2024enof), reaching new SOTA accuracies of 96.11% and 78.15%, respectively. Moreover, notably, as the time step increases, CKDSNN continues to significantly outperform prior works.
@@
SEW-R (fang2021sew) NIPS’21 4 63.18 67.04
RMP-Loss (guo2023rmp) ICCV’23 4 63.14 64.71
MBPN (guo2023mbpn) ICCV’23 4 63.03 65.17
@@
KDSNN (Xu2023kdsnn) CVPR’23 4 63.42 67.18
EnOFSNN (guo2024enof) NIPS’24 4 65.31 67.40
BKDSNN (xu2025bkdsnn) ECCV’24 4 65.60 71.24

### v2

@@ Introduction
Spiking Neural Networks (SNNs), inspired by the spiking mechanism of biological neurons, utilize event-driven binary spikes to transmit information, allowing multiplications between activations and weights to be replaced by additions or remain silent, thereby significantly improving energy efficiency (Eshraghian et al. 2023; Davies et al. 2018). Taking advantage of this computational paradigm, SNNs can operate efficiently on neuromorphic hardware and demonstrate autonomous learning capabilities and ultralow power consumption (Mehonic and Kenyon 2022), making them highly promising for intelligent computing tasks (Fang et al. 2023; Zhang et al. 2020). However, training SNNs presents significant challenges due to the inherently discrete and sparse nature of spike-based features, which complicates their optimization process and results in performance and application limitations compared to traditional artificial neural networks (ANNs) (Zhou et al. 2021, 2022; Zhou, Guo, and Wang 2023; Zhou et al. 2024d, a, b, e, c; Li et al. 2025, 2024; Qian et al. 2025; Zhao et al. 2025; Zhou et al. 2025a, b; Jin et al. 2025; Zhou et al. 2025c; Kryklyvets et al. 2025; Shen et al. 2023; Song et al. 2022; Guo et al. 2025).
@@
Specifically, conversion-based methods (Bu et al. 2022) transfer pre-trained ANN parameters to corresponding SNNs but replace ReLU activation with spiking neurons (Huang et al. 2024). This strategy has been experimentally found to require a large number of time steps to achieve satisfactory performance (Bu et al. 2022; Han, Srinivasan, and Roy 2020; Rueckauer et al. 2017). Direct training methods (Wu et al. 2019; Fang et al. 2021; Deng et al. 2022), on the other hand, optimize SNNs using direct backpropagation through the surrogate gradient estimation technique (Zenke and Vogels 2021), leading to significant progress in pattern recognition (Deng et al. 2022; Zhou et al. 2023), natural language processing  (Zhang et al. 2024; Xiao et al. 2022), and multimodal tasks (Liu et al. 2025). While this strategy reduces training time steps, the SNN’s performance still lags behind that of ANNs.
@@
Recent works (Xu et al. 2024, 2023) improve SNN training using knowledge distillation (KD) (Hinton, Vinyals, and Dean 2014) techniques, with pretrained ANNs as the teacher model and SNNs as the student model, showing promising results on multiple datasets. These methods perform KD by element-wise aligning the output features (Xu et al. 2023) or classification logits (Xu et al. 2023) between ANNs and SNNs. Such a KD paradigm distills knowledge from teacher ANNs to student SNNs, however, prior works ignore two critical issues: (i) Discrepancy between raw features. As illustrated in Fig. 1(a), features extracted from ANNs via a single forward pass are represented as continuous floating-point values. In contrast, features in SNNs, obtained by forward propagation over multiple time steps, are expressed as discrete binary spikes. Moreover, whereas ANN features encapsulate patterns spanning the entire image, SNN spikes primarily highlight salient regions. (ii) Discrepancy between raw logits. The logits (i.e., raw classification scores) are derived from the hidden features. Consequently, a notable disparity emerges between the raw logits of the two models. Specifically, SNN logits display greater sparsity and a more peaked distribution relative to those of ANNs.
@@
In this paper, we propose novel knowledge distillation strategies, having a Closer look at KD for SNN. Specifically, to address the first challenge, we propose the Saliency-scaled Activation Map Distillation (SAMD). As illustrated in Fig. 1, unlike previous methods that directly perform knowledge distillation using raw features, our SAMD leverages the Class Activation Map (CAM) (Zhou et al. 2016) of the teacher ANN, which provides more precise and focused knowledge, clearly describing the salient image regions related to the target class. Notably, unlike traditional activation map-based distillation methods from ANNs (e.g., e2KD (Parchami-Araghi et al. 2024) and CATKD (Guo et al. 2023c)), we discover that the surrogate gradient estimation in SNNs prevents the use of precise gradient estimation methods like Grad-CAM (Selvaraju et al. 2017) to generate saliency activation maps. Instead, we redesign the activation map distillation by aligning the Spiking Activation Map (SAM) of the student SNN with the CAM of the teacher ANN. Although both SAM and CAM originate from features, they are more consistent than raw features due to the use of saliency maps. Furthermore, we consider the numerical magnitude differences between the activation maps generated from SNN features (i.e., SAM) and ANN features (i.e., CAM). To more accurately assess the contribution of each pixel in the saliency maps, we apply the softmax function to convert both CAM and SAM into probability distributions. In this way, the scaled CAM and SAM remain consistent in both semantics and numerical magnitude, facilitating their alignment. To address the second challenge, we propose the Noise-smoothed Logits Distillation (NLD). As demonstrated in Fig. 1(b), NLD employs Gaussian noise to moderate the prediction logits of student SNNs. Specifically, we sample Gaussian noise with mean and variance parameters derived from the SNN logits, ensuring that the original distribution of these logits remains largely preserved. After the addition of noise, the logits of SNNs transit from a sparse and sharply peaked distribution to one that is denser and broader, resembling the distribution of ANN logits, thereby facilitating knowledge transmission between teacher and student. We evaluate the effectiveness and superiority of the proposed two KD strategies on CIFAR-10, CIFAR-100, and ImageNet-1K da of SNNs.
@@
Spiking Neural Networks (SNNs) are brain-inspired models that mimic biological neural systems by transmitting information through discrete spikes, achieving lower energy consumption. SNNs are typically trained using two main approaches: ANN-SNN conversion (Meng et al. 2022; Bu et al. 2022; Deng and Gu 2021; Hu, Tang, and Pan 2023) and direct training  (Fang et al. 2021; Guo et al. 2023a, b; Deng et al. 2022; Meng et al. 2023). The conversion methods directly transform pre-trained ANN into SNN by replacing its ReLU activation functions with integrate-and-fire (IF) (Bu et al. 2022) neurons. However, the converted SNNs often require prolonged time steps to collect sufficient spike signals to ensure accuracy. The direct training methods directly train the SNN by backpropagating surrogate gradients (Fang et al. 2021) through multiple time steps, which can alleviate excessive time steps, enabling efficient training and inference. However, a significant accuracy gap persists between the trained SNNs and ANNs due to the approximation errors (Xu et al. 2023) inherent in surrogate gradients. Unlike them, our method utilizes the informative knowledge in pretrained ANNs to better supervise SNN training, achieving balance between accuracy and efficiency.
@@
Knowledge Distillation (KD) (Hinton, Vinyals, and Dean 2014) was initially proposed for compressing artificial neural networks (ANNs) by transferring knowledge from complex teacher models to lightweight student models. Traditional logits-based KD methods typically minimize the difference in classification probabilities (i.e., logits) between the student and teacher models (Sun et al. 2024; Jin, Wang, and Lin 2023). In addition, feature-based KD methods enhance the performance of the student model by learning the feature representations of the teacher model (Wang et al. 2025; Guo et al. 2023c). Notably, a special class of feature-based KD methods (Guo et al. 2023c; Parchami-Araghi et al. 2024; Zagoruyko and Komodakis 2017) enhances the performance of the student by minimizing the semantic saliency activation maps of both the teacher and student models (e.g., Grad-CAM (Selvaraju et al. 2017)). Recently, a few works start to apply KD technique to facilitate SNN training (Xu et al. 2024, 2023), significantly improving the model’s performance. i.e., KDSNN (Xu et al. 2023) regularizes consistency of the output features and logits between ANNs and SNNs, while BKDSNN (Xu et al. 2024) enhances feature-level matching by further processing the spike features of SNNs with a blurring matrix. However, these methods largely overlook the discrepancies of the raw features and logits from teacher ANNs and student SNNs. Specifically, ANNs generate continuous floating-point features, while SNNs generate discrete spike features. The logits of SNNs are also more sparse than that of ANNs. Thus, we perform KD from the perspective of semantic saliency activation maps, which are more semantically aligned. But, prior activation map-based KD methods (Parchami-Araghi et al. 2024; Zagoruyko and Komodakis 2017) from ANNs cannot be directly applied to SNNs, due to the surrogate gradient estimation errors in SNNs, which affect the generation of semantic saliency activation maps. Besides, CATKD (Guo et al. 2023c) is only applicable to CNN-based architectures, which limits the application of semantic saliency activation map-based KD in SNNs. In contrast, we design a semantic saliency activation map-based KD method for the characteristics of SNNs, which can be applied to various architectures. Moreover, we also design a noise-smoothing distillation method at the logits-level to further enhance the performance of SNNs.
@@
The overall pipeline of our CKDSNN framework is illustrated in Fig. 2. Given a pretrained ANN teacher model and a learnable SNN student model, the proposed CKDSNN aims to train the student model by effectively distilling knowledge from the teacher model from two aspects: 1) Saliency-scaled Activation Map Distillation (SAMD). The class activation map (Selvaraju et al. 2017) obtained from the teacher ANN is used to guide the student SNN to fire spikes in salient image regions. During the distillation process, we address the magnitude discrepancy between the two types of activation maps by scaling them into the same range. 2) Noise-smoothed Logits Distillation (NLD). The classification output logits of the student model is smoothed using additional Gaussian noise, making the vanilla sparse logits distribution of student SNN close to the continuous logits distribution of teacher ANN. This facilitates more precise logit-level knowledge distillation. We first have a brief introduction to the spiking neuron used in SNN, which leads to the discrepancy on the features and logits of ANN and SNN. Then, we elaborate on the proposed SAMD and NLD strategies, respectively.
@@
Typical SNNs (Fang et al. 2021; Deng et al. 2022; Meng et al. 2023; Jiang et al. 2024; Deng et al. 2024) adopt the integrate-and-fire (IF) neuron as the fundamental unit. Specifically, the IF neuron first integrates input currents by updating its membrane potential, and then compares it with a pre-set threshold to generate a spike signal, followed by a reset mechanism of the membrane potential. This process can be formulated as:
@@
where $I[t]$ is the input current at time step $t$ and V[t1]V[t{-}1] is the membrane potential at previous t1t{-}1 time step. $\gamma$ denotes the membrane time constant and $\Theta(\cdot)$ represents the Heaviside function (Huang et al. 2024). The IF accumulates input currents to update the membrane potential $H[t]$. When $H[t]$ exceeds the threshold $V_{\text{th}}$, a spike $S[t]$ is generated, and the membrane potential is reset to $V_{\text{reset}}$ (Meng et al. 2023).
In image classification, the above is attached to each feature encoding block of SNN. Specifically, each encoding block contains multiple linear transformation layers, batch normalization layers, and IF spiking neurons, as shown in Fig. 2(a). This generates discrete spike features $\boldsymbol{F}^{st}{\in}\mathbb{R}^{T\times C\times H\times W}$ over $T$ time steps, where $C$, $H$, and $W$ denote the number of channels, height, and width of the feature, respectively. The discrete spike features $\boldsymbol{F}^{st}$ generated by the IF activation mechanism differ significantly from the continuous features $\boldsymbol{F}^{te}{\in}\mathbb{R}^{C\times H\times W}$ produced by ANN. This leads to a significant discrepancy in the feature distribution between $\boldsymbol{F}^{st}$ and $\boldsymbol{F}^{te}$, resulting in different prediction logits. Prior KD methods (Xu et al. 2023) directly match these two types of features or logits element-wise, ignoring the essential differences in feature representation and distribution, leading to suboptimal distillation results. We address these issues through carefully designed KD strategies.
@@
We propose the Saliency-scaled Activation Map Distillation (SAMD) to train SNN rather than distilling the element-wise discrepant features used in existing works (Xu et al. 2023, 2024). Specifically, SAMD consists of three steps:
1) Class Activation Map (CAM) Generation. The class activation map is initially used as a visualization tool to enhance the explainability of convolutional neural network (Zhou et al. 2016), which can highlight the related image region of a specific class. In our work, we use it as the teacher’s knowledge for distillation. We follow the typical method of Grad-CAM (Selvaraju et al. 2017) to generate the class activation map. Specifically, given an input image $x$, we first extract the intermediate features using a pretrained teacher ANN model, denoted as $\boldsymbol{F}^{te}{\in}\mathbb{R}^{C\times H\times W}$. Then, the gradients of $\boldsymbol{F}^{te}$ are calculated according to the forward loss between prediction and the ground truth label $y{\in}\mathbb{R}^{K}$ ($K$ is the total number of classes). The class activation map $\boldsymbol{M}^{te}{\in}\mathbb{R}^{H\times W}$ can be obtained by associating $\boldsymbol{F}^{te}$ with the class label $y$. This process can be formulated as follows:
@@
In addition to the feature-driven activation map distillation, we consider the target-driven logits distillation, which aims to align the output probability logits of teacher ANN and student SNN. Although prior work (Xu et al. 2023) considered this, they largely overlook the discrepancy between the logits. This issue still originates from the utilization of different features in logits production where SNNs employ binary spike features that are either 0 or 1, whereas ANNs utilize floating-point features. Consequently, the logits of SNN are very sparse, while those of ANN are more dense and continuous. We propose the Noise-Smoothed Logits Distillation (NLD) to address this problem. Thus, we try to soften the logits of student SNN, $z^{st}{\in}\mathbb{R}^{1\times K}$, reducing the discrepancy from $z^{te}{\in}\mathbb{R}^{1\times K}$. Specifically, we add some continuous noise $\epsilon{\in}\mathbb{R}^{1\times K}$ onto $z^{st}$. To avoid destroying the original distribution of $z^{st}$ while maintaining the characteristic of classification logits, we opt to Gaussian noise with the mean and standard deviation of $z^{st}$:
@@
Datasets. We evaluate our method on three widely used image classification datasets and a neuromorphic dataset in this research area: CIFAR-10/100 (Krizhevsky and Hinton 2009), ImageNet-1K (Deng et al. 2009) and CIFAR10-DVS (Li et al. 2017). The details of these datasets are provided in the supplementary materials.
Model Configuration. To ensure fair comparison, we determine the configurations of teacher ANN and student SNN models based on prior studies (Guo et al. 2024; Xu et al. 2024). Specifically, for the CIFAR-10/100 datasets, we use the ResNet-19/20 versions provided by TET (Deng et al. 2022) and QCFS (Li et al. 2023), with their corresponding ANN versions as the teacher models; and we also test a ViT-based architecture, using ViT-S as the teacher ANN and Spikformer-4-384 (Zhou et al. 2023) as the student model. For the ImageNet-1K dataset, we use the ResNet-18 and ResNet-34 pre-trained on ImageNet-1K as teacher models, while SEW-ResNet (Fang et al. 2021) serves as the student model. For the CIFAR10-DVS dataset, we use ResNet-19 as the student, with its ANN architecture serving as the teacher, trained in the same way as EnOF (Guo et al. 2024).
Implementation Details. To ensure consistency with prior studies, our implementation on CIFAR-10/100 and ImageNet strictly follows the established distillation architecture (Guo et al. 2024; Xu et al. 2024). Specifically, we set the hyperparameters as follows: $\mathcal{T}$ in Eq. Saliency-scaled Activation Map Distillation and $\tau$ in Eq. 7 to $2.0$. $\lambda$ in Eq. 6 to $0.1$, $\beta$ and $\gamma$ in Eq. 8 to $1.0$. The Integrate-and-Fire (IF) neuron settings align with prior works (Fang et al. 2021), and all other training configurations, including batch size, learning rate, and optimizer, remain consistent with those in (Xu et al. 2024). In addition, we conduct sensitivity analysis of all hyper-parameters in the supplementary materials.
Platform. All experiments are conducted on a server platform equipped with 32 cores Intel Xeon Platinum 8352V CPU with 2.10GHz and 8-way NVIDIA GPUs. We use SpikingJelly (Fang et al. 2023) to simulate the IF (Tal and Schwartz 1997) spiking neurons.
@@
SEW-R (Fang et al. 2021) NIPS’21 4 89.07 60.16 93.24 70.84 - -
STBP (Wu et al. 2019) AAAI’21 4 - - 92.92 - - -
TET (Deng et al. 2022) ICLR’22 4 - - 94.44 74.47 - -
SLTT (Meng et al. 2023) ICCV’23 4 - - 94.56 74.67 - -
Spikformer (Zhou et al. 2023) ICLR’22 4 - - - - 95.93 79.65
@@
KDSNN (Xu et al. 2023) CVPR’23 4 89.03 60.18 94.36 74.08 95.88 80.33
BKDSNN (Xu et al. 2024) ECCV’24 4 89.29 60.92 94.64 74.95 96.06 81.26
EnOFSNN (Guo et al. 2024) NIPS’24 1 92.66 70.38 95.37 77.08 - -
@@
Results on CIFAR-10/100. The comparison results between our CKDSNN and previous methods are shown in Tab. 1. We significantly outperform existing methods across different architectures. For example, in the ResNet-19, when the time step is set to 1, CKDSNN achieves an accuracy improvement of 0.74% on the CIFAR-10 dataset and 1.03% on the CIFAR-100 dataset compared to the current most competitive knowledge distillation method, EnOF (Guo et al. 2024), reaching new SOTA accuracies of 96.11% and 78.15%, respectively. Moreover, notably, as the time step increases, CKDSNN continues to significantly outperform prior works.
@@
SEW-R (Fang et al. 2021) NIPS’21 4 63.18 67.04
RMP-Loss (Guo et al. 2023a) ICCV’23 4 63.14 64.71
MBPN (Guo et al. 2023b) ICCV’23 4 63.03 65.17
@@
KDSNN (Xu et al. 2023) CVPR’23 4 63.42 67.18
EnOFSNN (Guo et al. 2024) NIPS’24 4 65.31 67.40
BKDSNN (Xu et al. 2024) ECCV’24 4 65.60 71.24